# Kernelized Concept Erasure

### v5

@@ A.2 Kernelization of the Minimax Game
The projection onto the orthogonal complement of $\boldsymbol{w}=\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})$ is defined as the following
$\mathrm{P}_{\boldsymbol{w}}^{\perp}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\mathrm{I}-\frac{\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{i})}\Big{)}\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})^{\top}}\Big{)}}{\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})^{\top}\Big{)}\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})\Big{)}}$ (17)
@@
$\mathrm{P}_{\boldsymbol{w}}^{\perp}\boldsymbol{\Phi}(\mathbf{z})$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\mathrm{I}\boldsymbol{\Phi}(\mathbf{z})-\frac{\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})}\Big{)}\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n^{\prime}})^{\top}}\Big{)}}{\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})^{\top}\Big{)}\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})\Big{)}}\,\boldsymbol{\Phi}(\mathbf{z})$ (18a)
$=\boldsymbol{\Phi}(\mathbf{z})-\frac{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\boldsymbol{\Phi}(\mathbf{x}_{n})\boldsymbol{\Phi}(\mathbf{x}_{m})^{\top}}\boldsymbol{\Phi}(\mathbf{z})}{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\boldsymbol{\Phi}(\mathbf{x}_{n})^{\top}\boldsymbol{\Phi}(\mathbf{x}_{m})}}$ (18b)
$=\boldsymbol{\Phi}(\mathbf{z})-\frac{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\boldsymbol{\Phi}(\mathbf{x}_{n})\kappa(\mathbf{x}_{m},\mathbf{z})}}{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\kappa(\mathbf{x}_{n},\mathbf{x}_{m})}}$ (18c)
$=\boldsymbol{\Phi}(\mathbf{z})-\underbrace{\left(\frac{\sum_{m=1}^{N}{\alpha_{m}\kappa(\mathbf{x}_{m},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)}_{\in\mathbb{R}}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (18d)
$=\boldsymbol{\Phi}(\mathbf{z})-\left(\frac{\sum_{m=1}^{N}{\alpha_{m}\kappa(\mathbf{x}_{n},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\boldsymbol{w}$ (18e)
=$\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\boldsymbol{\Phi}_{\text{proj}}(\mathbf{z})}$ (18f)
@@
${\boldsymbol{\theta}}=\sum_{n=1}^{N}\beta_{n}\boldsymbol{\Phi}(\mathbf{x}_{n})$ (19)
@@
$\Big{\langle}{\boldsymbol{\theta}},$ ${\boldsymbol{\Phi}_{\text{proj}}(\mathbf{z})}\Big{\rangle}=\Big{\langle}{\boldsymbol{\theta}},\boldsymbol{\Phi}(\mathbf{z})-\left(\frac{\sum_{n=1}^{N}{\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\boldsymbol{w}\Big{\rangle}$ (20a)
@@
$=\Big{\langle}\sum_{m=1}^{N}\beta_{m}\boldsymbol{\Phi}(\mathbf{x}_{m}),\boldsymbol{\Phi}(\mathbf{z})-\left(\frac{\sum_{n=1}^{N}{\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\boldsymbol{w}\Big{\rangle}$ (20b)
@@
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\left(\frac{\sum_{n=1}^{N}{\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\kappa(\mathbf{x}_{m},\boldsymbol{w})$ (20c)
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\frac{1}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\sum_{n=1}^{N}\sum_{m=1}^{N}\alpha_{n}\beta_{m}\kappa(\mathbf{x}_{n},\mathbf{z})\kappa(\mathbf{x}_{m},\boldsymbol{w})$ (20d)
@@
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})\kappa(\mathbf{x}_{m},\boldsymbol{w})\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20e)
@@
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})\boldsymbol{\Phi}(\mathbf{x}_{m})^{\top}\left(\sum_{n^{\prime}=1}^{N}\alpha_{n^{\prime}}\boldsymbol{\Phi}(\mathbf{x}_{n^{\prime}})\right)\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20f)
@@
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\sum_{n^{\prime}=1}^{N}\alpha_{n^{\prime}}\alpha_{n}\kappa(\mathbf{x}_{n},\mathbf{z})\boldsymbol{\Phi}(\mathbf{x}_{m})^{\top}\boldsymbol{\Phi}(\mathbf{x}_{n^{\prime}})\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20g)
@@
$=\sum_{m=1}^{N}\beta_{m}\kappa(\mathbf{x}_{m},\mathbf{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\sum_{n^{\prime}=1}^{N}\alpha_{n}\alpha_{n^{\prime}}\kappa(\mathbf{x}_{n},\mathbf{z})\kappa(\mathbf{x}_{m},\mathbf{x}_{n^{\prime}})\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20h)
@@
$=\sum_{m=1}^{N}\beta_{m}\left(\kappa(\mathbf{x}_{m},\mathbf{z})-\frac{\boldsymbol{\alpha}^{\top}K^{(m)}(\mathbf{z})\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)$ (20i)
@@
where we define the following matrix component-wise: $K^{(m)}_{ij}(\mathbf{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\kappa(\mathbf{x}_{i},\mathbf{z})\kappa(\mathbf{x}_{m},\mathbf{x}_{j})$. Eq. 20i can be evaluated without explicitly applying the kernel transformation $\boldsymbol{\Phi}$. In terms of notation, when we have $\Big{\langle}{\boldsymbol{\theta}},{\boldsymbol{\Phi}_{\text{proj}}}(\mathbf{z}_{n})\Big{\rangle}$, we write
@@
$\sum_{m=1}^{N}\beta_{m}\left(\kappa(\mathbf{x}_{m},\mathbf{z}_{n})-\frac{\boldsymbol{\alpha}^{\top}K^{(m,n)}\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)$ (21)
where we define $K^{(m,n)}_{ij}(\mathbf{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\kappa(\mathbf{x}_{i},\mathbf{z}_{n})\kappa(\mathbf{x}_{m},\mathbf{x}_{j})$. This proves the result. ∎

### v6

@@ A.2 Kernelization of the Maximin Game
The projection onto the orthogonal complement of $\boldsymbol{w}=\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})$ is defined as the following
$\mathrm{P}_{\boldsymbol{w}}^{\perp}\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\mathrm{I}-\frac{\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\mathbf{x}_{i})}\Big{)}\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})^{\top}}\Big{)}}{\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})^{\top}\Big{)}\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})\Big{)}}$ (17)
@@
$\mathrm{P}_{\boldsymbol{w}}^{\perp}\boldsymbol{\Phi}(\boldsymbol{z})$ =$\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}\mathrm{I}\boldsymbol{\Phi}(\boldsymbol{z})-\frac{\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})}\Big{)}\Big{(}\sum_{n=1}^{N}{\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n^{\prime}})^{\top}}\Big{)}}{\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})^{\top}\Big{)}\Big{(}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})\Big{)}}\,\boldsymbol{\Phi}(\boldsymbol{z})$ (18a)
$=\boldsymbol{\Phi}(\boldsymbol{z})-\frac{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\boldsymbol{\Phi}(\boldsymbol{x}_{n})\boldsymbol{\Phi}(\boldsymbol{x}_{m})^{\top}}\boldsymbol{\Phi}(\boldsymbol{z})}{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\boldsymbol{\Phi}(\boldsymbol{x}_{n})^{\top}\boldsymbol{\Phi}(\boldsymbol{x}_{m})}}$ (18b)
$=\boldsymbol{\Phi}(\boldsymbol{z})-\frac{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}\boldsymbol{\Phi}(\boldsymbol{x}_{n}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})}}{\sum_{n=1}^{N}\sum_{m=1}^{N}{\alpha_{n}\alpha_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{x}_{m})}}$ (18c)
$=\boldsymbol{\Phi}(\boldsymbol{z})-\underbrace{\left(\frac{\sum_{m=1}^{N}{\alpha_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)}_{\in\mathbb{R}}\sum_{n=1}^{N}\alpha_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})$ (18d)
$=\boldsymbol{\Phi}(\boldsymbol{z})-\left(\frac{\sum_{m=1}^{N}{\alpha_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\boldsymbol{w}$ (18e)
=$\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\boldsymbol{\Phi}_{\text{proj}}(\boldsymbol{z})}$ (18f)
@@
${\boldsymbol{\theta}}=\sum_{n=1}^{N}\beta_{n}\boldsymbol{\Phi}(\boldsymbol{x}_{n})$ (19)
@@
$\Big{\langle}{\boldsymbol{\theta}},$ ${\boldsymbol{\Phi}_{\text{proj}}(\boldsymbol{z})}\Big{\rangle}=\Big{\langle}{\boldsymbol{\theta}},\boldsymbol{\Phi}(\boldsymbol{z})-\left(\frac{\sum_{n=1}^{N}{\alpha_{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\boldsymbol{w}\Big{\rangle}$ (20a)
@@
$=\Big{\langle}\sum_{m=1}^{N}\beta_{m}\boldsymbol{\Phi}(\boldsymbol{x}_{m}),\boldsymbol{\Phi}(\boldsymbol{z})-\left(\frac{\sum_{n=1}^{N}{\alpha_{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)\boldsymbol{w}\Big{\rangle}$ (20b)
@@
$=\sum_{m=1}^{N}\beta_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})-\sum_{m=1}^{N}\beta_{m}\left(\frac{\sum_{n=1}^{N}{\alpha_{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{z})}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{w})$ (20c)
$=\sum_{m=1}^{N}\beta_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})-\frac{1}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\sum_{n=1}^{N}\sum_{m=1}^{N}\alpha_{n}\beta_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{z}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{w})$ (20d)
@@
$=\sum_{m=1}^{N}\beta_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\alpha_{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{z}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{w})\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20e)
@@
$=\sum_{m=1}^{N}\beta_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\alpha_{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{z})\boldsymbol{\Phi}(\boldsymbol{x}_{m})^{\top}\left(\sum_{n^{\prime}=1}^{N}\alpha_{n^{\prime}}\boldsymbol{\Phi}(\boldsymbol{x}_{n^{\prime}})\right)\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20f)
@@
$=\sum_{m=1}^{N}\beta_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\sum_{n^{\prime}=1}^{N}\alpha_{n^{\prime}}\alpha_{n}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{z})\boldsymbol{\Phi}(\boldsymbol{x}_{m})^{\top}\boldsymbol{\Phi}(\boldsymbol{x}_{n^{\prime}})\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20g)
@@
$=\sum_{m=1}^{N}\beta_{m}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})-\sum_{m=1}^{N}\beta_{m}\frac{\left(\sum_{n=1}^{N}\sum_{n^{\prime}=1}^{N}\alpha_{n}\alpha_{n^{\prime}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{n},\boldsymbol{z}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{x}_{n^{\prime}})\right)}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}$ (20h)
@@
$=\sum_{m=1}^{N}\beta_{m}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z})-\frac{\boldsymbol{\alpha}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}^{(m)}}(\boldsymbol{z})\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}}\boldsymbol{\alpha}}\right)$ (20i)
@@
where we define the following matrix component-wise: $\boldsymbol{K}^{(m)}_{ij}(\boldsymbol{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{i},\boldsymbol{z}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{x}_{j})$. Eq. 20i can be evaluated without explicitly applying the kernel transformation $\boldsymbol{\Phi}$. In terms of notation, when we have $\Big{\langle}{\boldsymbol{\theta}},{\boldsymbol{\Phi}_{\text{proj}}}(\boldsymbol{z}_{n})\Big{\rangle}$, we write
@@
$\sum_{m=1}^{N}\beta_{m}\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{z}_{n})-\frac{\boldsymbol{\alpha}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{K}^{(m,n)}}\boldsymbol{\alpha}}{\boldsymbol{\alpha}^{\top}K\boldsymbol{\alpha}}\right)$ (21)
where we define $\boldsymbol{K}^{(m,n)}_{ij}(\boldsymbol{z})\mathrel{\stackrel{{\scriptstyle\textnormal{def}}}{{=}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{i},\boldsymbol{z}_{n}){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\kappa}(\boldsymbol{x}_{m},\boldsymbol{x}_{j})$.author=ryan,color=violet!40,size=,fancyline,caption=,]should $\boldsymbol{z}_{n}$ be $\boldsymbol{z}$? This proves the result. ∎